Operations Report – September 28th

MDRS Operations Report 25-28-SEPT-2021

Name of person filing report: Shannon Rupert and Crew Engineer David Laude

Reason for Report: Routine daily

Non-nominal systems: Autoclave

Action taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Off

Solar— Nominal, providing all power

Solar— SOC % Last 24 hours:

Average 68.3

Minimum 47

Maximum 100

Note on solar: Nothing to report

Diesel Reading – Empty, tank to be removed soon

Propane Reading, station tank – 80%

Propane Reading, director tank— 74%

Propane Reading, intern tank— 72%

Propane Reading, generator— 80%

Ethanol Free Gasoline – 5 gallons

Water (loft tank): 41 gallons

Water Meter: 152,297.4 gallons

Water (static tank) – ~ 450 gallons

Static to Loft Pump used – yes

Water in GreenHab – gallons Unknown to me at this time. Will report after I find out how to measure it. Too late now.

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 130 gallons

Hab toilet tank emptied: no

Perseverance rover used: yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Sojourner rover used: yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Curiosity rover used: yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: We will get hours on the first EVA.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): Yes

Reason for use: parked at Outpost

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 1

Notes on ATVs: One of the 350’s is still acting as a sign because the battery is dead. I will move tomorrow

HabCar used and why, where? Yes, for water and supplies

CrewCar used and why, where? Yes, to Grand Junction to pick up crew

Luna used and why, where? Yes, but only moved on campus

General notes and comments: Nothing to report

Summary of internet: All old accounts are nominal. New service still not working.

Summary of suits and radios: We tried on both styles. All of us picked the single piece unit. The two piece is too flawed. The lower helmet ring is too far forward as it causes the head to be forced against the back of head restricting movement and constant muscle tension against it. Also, the rear ring clamps are forced against the frame making them difficult to open or close. On some helmets the upper ring front down-pointing latch section is bent backward causing it not to engage lower clamp mate. That is easily fixed with pliers, but every time the helmet is set down the front rests upon that small piece. We would send photos if requested. Regarding the one piece, unit 7 had no fan when switched on. One battery connector had slipped off a battery terminal. Reattachment was firm, but since it came off, I will make it hold better tomorrow by crimping the female connector. One radio has suddenly become erratic and unusable. With an in-situ crew of 4, we have spares.

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: RAM and SciDome were cleaned. Four signs built by the work party had pier blocks attached and they were deployed-two along Cow Dung Road stating where our land begins and ends, and two on the driveways to restrict unauthorized personnel from driving to the station. So far tourist reactions are mixed.

Summary of Hab operations: All but a few food items supplied to the crew. Lower deck cleaning completed.

Summary of Outpost operations: Nothing to report.

Summary of GreenHab operations:

Summary of ScienceDome operations: Building was cleaned today but still some things to do before it is complete

Summary of RAM operations: Organized and cleaned. Tools used to assemble 4 heavy warning signs.

Summary of any observatory issues: Robotic observatory is down.

Summary of health and safety issues: All crew nominal.

Questions, concerns, supplies needed and requests: A few food items to come. Body temp thermometer coming. Smoke detector for RAM requested.